# https://tet.uni-wuppertal.de

### v1

@@ 3 Stabilized EMQS Formulations with Auxiliary Variables
$\begin{split}\underbrace{\left[\begin{array}[]{cccc}0&\varepsilon\mathrm{grad}&0&\varepsilon\mathrm{grad}\\ \mathrm{div}\varepsilon&\mathrm{div}\varepsilon\mathrm{grad}&0&0\\ 0&0&\mu&0\\ \mathrm{div}\varepsilon&0&0&0\end{array}\right]}_{{\mathbf{E}}_{\mathrm{EMQS3}}:=}\frac{\partial}{\partial t}\left[\begin{array}[]{c}\!\frac{\partial}{\partial t}\bm{A}\!\\ \!\varphi\!\\ \!{\bm{H}}\!\\ \!\lambda\!\end{array}\right]\!\!=\!\!\ \left[\left[\begin{array}[]{cccc}0&0&-\mathrm{curl}&0\\ 0&0&0&0\\ \mathrm{curl}&0&0&0\\ 0&0&0&0\end{array}\right]\right.\\ -\left.\left[\begin{array}[]{cccc}\kappa&\kappa\mathrm{grad}&0&0\\ \mathrm{div}\kappa&\mathrm{div}\kappa\mathrm{grad}&0&0\\ 0&0&0&0\\ 0&0&0&0\end{array}\right]\right]\left[\begin{array}[]{c}\!\frac{\partial}{\partial t}\bm{A}\!\\ \!\varphi\!\\ \!{\bm{H}}\!\\ \!\lambda\!\end{array}\right]\!\!+\!\!\left[\begin{array}[]{c}\!1\!\\ \!\mathrm{div}\!\\ \!0\!\\ \!0\!\end{array}\right]\bm{j}_{\mathrm{s}},\end{split}$ (13)
with Lagrange multiplier $\lambda$, input $\mathrm{u}(t)=\bm{j}_{\mathrm{s}}$ and $\mathrm{y}(t)=\left(\frac{\partial}{\partial t}\bm{A}+\mathrm{grad}\varphi\right)=-\bm{E}.$ System (13) is a PHS, where the compatibility relation (4) holds for the Hamiltonian (12).
The Lagrange multiplier $\lambda$ corresponds to a second scalar electric potential and as solutions to (13) exists, where $\mathrm{div}\left(\varepsilon\frac{\partial^{2}}{\partial t^{2}}\bm{A}\right)=0,$ holds and $\lambda=0,$ it is possible to modify (13) to a formulation that corresponds to a split potential formulation for the electric field $\bm{E}=-\frac{\partial}{\partial t}\bm{A}-\mathrm{grad}(\varphi+\lambda)$ as introduced in [5] via
$\begin{split}\underbrace{\left[\begin{array}[]{cccc}0&\varepsilon\mathrm{grad}&0&\varepsilon\mathrm{grad}\\ \mathrm{div}\varepsilon&\mathrm{div}\varepsilon\mathrm{grad}&0&\mathrm{div}\varepsilon\mathrm{grad}\\ 0&0&\mu&0\\ \mathrm{div}\varepsilon&\mathrm{div}\varepsilon\mathrm{grad}&0&\mathrm{div}\varepsilon\mathrm{grad}\end{array}\right]}_{{\mathbf{E}}_{\mathrm{EMQS4}}:=}\frac{\partial}{\partial t}\left[\begin{array}[]{c}\!\frac{\partial}{\partial t}\bm{A}\!\\ \!\varphi\!\\ \!\bm{H}\!\\ \!\lambda\!\end{array}\right]\!\!=\!\!\ \left[\left[\begin{array}[]{cccc}0&0&-\mathrm{curl}&0\\ 0&0&0&0\\ \mathrm{curl}&0&0&0\\ 0&0&0&0\end{array}\right]\right.\\ -\left.\left[\begin{array}[]{cccc}\kappa&\kappa\mathrm{grad}&0&0\\ \mathrm{div}\kappa&\mathrm{div}\kappa\mathrm{grad}&0&0\\ 0&0&0&0\\ 0&0&0&0\end{array}\right]\right]\left[\begin{array}[]{c}\!\frac{\partial}{\partial t}\bm{A}\!\\ \!\varphi\!\\ \!\bm{H}\!\\ \!\lambda\!\end{array}\right]\!\!+\!\!\left[\begin{array}[]{c}\!1\!\\ \!\mathrm{div}\!\\ \!0\!\\ \!\mathrm{div}\!\end{array}\right]\bm{j}_{\mathrm{s}},\end{split}$ (14)
@@
$H_{\mathrm{EMQS}}=\frac{1}{2}\left(\varepsilon\|\mathrm{grad}\varphi\|^{2}+\varepsilon\|\mathrm{grad}\lambda\|^{2}+\mu\|\bm{H}\|^{2}\right)\\ +\varepsilon\frac{\partial}{\partial t}\bm{A}\cdot\mathrm{grad}(\varphi+\lambda)+\varepsilon\mathrm{grad}\varphi\cdot\mathrm{grad}\lambda$ (15)
corresponding the EMQS energy density, (4) holds for formulation (14) with
$\mathrm{grad}\,H_{\mathrm{EMQS}}=\mathbf{E}_{\mathrm{EMQS4}}^{\star}{\mathbf{z}}=\left[\begin{array}[]{c}\varepsilon\mathrm{grad}(\varphi+\lambda),\\ \mathrm{div}\left(\varepsilon(\frac{\partial}{\partial t}\bm{A}+\mathrm{grad}(\varphi+\lambda))\right)\\ \mu\bm{H}\\ \mathrm{div}\left(\varepsilon(\frac{\partial}{\partial t}\bm{A}+\mathrm{grad}(\varphi+\lambda))\right)\end{array}\right].$
@@
Symmetrized EMQS field formulations can also be set up to explicitly enforce the Coulomb-type gauge condition (8) via $\mathrm{div}\left(\hat{\kappa}\frac{\partial}{\partial t}\bm{A}\right)=\frac{1}{\tau}\mathrm{div}\left(\varepsilon\frac{\partial}{\partial t}\bm{A}\right)=0,$ with an artificial $\hat{\kappa}:=\frac{1}{\tau}\varepsilon$ (for a time constant $\tau$ to get the correct units) via Lagrange multipliers as e.g. in [12] reformulate as
$\begin{split}\underbrace{\left[\begin{array}[]{cccc}0&\varepsilon\mathrm{grad}&0&0\\ \mathrm{div}\varepsilon&\mathrm{div}\varepsilon\mathrm{grad}&0&0\\ 0&0&\mu&0\\ 0&0&0&0\end{array}\right]}_{{\mathbf{E}}_{\mathrm{EMQS4}}:=}\frac{\partial}{\partial t}\left[\begin{array}[]{c}\!\frac{\partial}{\partial t}\bm{A}\!\\ \!\varphi\!\\ \!\bm{H}\!\\ \!\lambda\!\end{array}\right]\!\!=\!\!\ \left[\left[\begin{array}[]{cccc}0&0&-\mathrm{curl}&0\\ 0&0&0&0\\ \mathrm{curl}&0&0&0\\ 0&0&0&0\end{array}\right]\right.\\ -\left.\left[\begin{array}[]{cccc}\kappa&\kappa\mathrm{grad}&0&\hat{\kappa}\mathrm{grad}\\ \mathrm{div}\kappa&\mathrm{div}\kappa\mathrm{grad}&0&0\\ 0&0&0&0\\ \mathrm{div}\hat{\kappa}&0&0&0\end{array}\right]\right]\left[\begin{array}[]{c}\!\frac{\partial}{\partial t}\bm{A}\!\\ \!\varphi\!\\ \!\bm{H}\!\\ \!\lambda\!\end{array}\right]\!\!+\!\!\left[\begin{array}[]{c}\!1\!\\ \!\mathrm{div}\!\\ \!0\!\\ \!0\!\end{array}\right]\bm{j}_{\mathrm{s}},\end{split}$ (17)
with the Lagrange multiplier $\lambda$, input $\mathrm{u}(t)=\bm{j}_{\mathrm{s}}$ and $\mathrm{y}(t)=\left(\frac{\partial}{\partial t}\bm{A}+\mathrm{grad}\varphi\right)=-\bm{E}.$ Note that $\hat{\kappa}$ can also be defined domain dependent, i.e., the gauge (8) can be defined in the whole computational domain or just in the non-conducting domains with the original electric conductivity $\kappa=0.$ The system (17) is a PHS, where the compatibility relation (4) holds for the Hamiltonian (12). With the explicitly enforced Coulomb-type gauge (8), the Hamiltonian (12) of the symmetric formulation (13) reduces to the expression
$H_{\mathrm{EMQS2}}=\frac{1}{2}\left(\varepsilon\|\mathrm{grad}\varphi\|^{2}+\mu\|\bm{H}\|^{2}\right).$ (18)
This expression corresponds the energy density of the EMQS field approximation, i.e., a formulation with the combined magnetic field and electro-quasistatic field energy.
@@ 4 Conclusions
Various infinite dimensional EMQS formulations have been classified within the port-Hamiltonian system framework, thus implying their numerical stability after a proper mimetic discretization where the PHS structure is found. The validity of the compatibility condition could be shown for symmetrized formulations, i.e., those EMQS formulations based on the Darwin-Ampere and the full Maxwell continuity equation. However, this re-introduced the second order time derivative of the magnetic vector potential that was discareded in the Darwin-Ampere equation under the no radiation EMQS assumption. For explicitely gauged EMQS formulations the relation between Lagrange multiplier and split scalar potential formulations is shown.

### v2

@@ 3 Stabilized EMQS Formulations with Auxiliary Variables
$\begin{split}\underbrace{\left[\begin{array}[]{cccc}0&\varepsilon\mathrm{grad}&0&\varepsilon\mathrm{grad}\\ -\color[rgb]{0,0,0}\mathrm{div}\varepsilon&-\color[rgb]{0,0,0}\mathrm{div}\varepsilon\mathrm{grad}&0&0\\ 0&0&\mu&0\\ -\color[rgb]{0,0,0}\mathrm{div}\varepsilon&0&0&0\end{array}\right]}_{{\mathbf{E}}_{\mathrm{EMQS3}}}\frac{\partial}{\partial t}\left[\begin{array}[]{c}\!\frac{\partial}{\partial t}\bm{A}\!\\ \!\varphi\!\\ \!{\bm{H}}\!\\ \!\lambda\!\end{array}\right]\!\!=\!\!\ \left[\left[\begin{array}[]{cccc}0&0&-\mathrm{curl}&0\\ 0&0&0&0\\ \mathrm{curl}&0&0&0\\ 0&0&0&0\end{array}\right]\right.\\ -\left.\left[\begin{array}[]{cccc}\kappa&\kappa\mathrm{grad}&0&0\\ -\color[rgb]{0,0,0}\mathrm{div}\kappa&-\color[rgb]{0,0,0}\mathrm{div}\kappa\mathrm{grad}&0&0\\ 0&0&0&0\\ 0&0&0&0\end{array}\right]\right]\left[\begin{array}[]{c}\!\frac{\partial}{\partial t}\bm{A}\!\\ \!\varphi\!\\ \!{\bm{H}}\!\\ \!\lambda\!\end{array}\right]\!\!+\!\!\left[\begin{array}[]{c}\!1\!\\ \!-\color[rgb]{0,0,0}\mathrm{div}\!\\ \!0\!\\ \!0\!\end{array}\right]\bm{j}_{\mathrm{s}},\end{split}$ (15)
with Lagrange multiplier $\lambda$, input $\mathrm{u}(t)=\bm{j}_{\mathrm{s}}$ and $\mathrm{y}(t)=\left(\frac{\partial}{\partial t}\bm{A}+\mathrm{grad}\varphi\right)=-\bm{E}.$ System (15) is a PHS, where the compatibility relation (4) holds for the Hamiltonian (14). The Lagrange multiplier $\lambda$ corresponds to a second scalar electric potential and as solutions to (15) exists, where $\mathrm{div}\left(\varepsilon\frac{\partial^{2}}{\partial t^{2}}\bm{A}\right)=0,$ holds and $\lambda=0,$ it is possible to modify (15) to a formulation that corresponds to a split potential formulation for the electric field $\bm{E}=-\frac{\partial}{\partial t}\bm{A}-\mathrm{grad}(\varphi+\lambda)$ as introduced in [9] via
$\begin{split}\underbrace{\left[\begin{array}[]{cccc}0&\varepsilon\mathrm{grad}&0&\varepsilon\mathrm{grad}\\ \color[rgb]{0.0,0.0,0.0}-\mathrm{div}\varepsilon&\color[rgb]{0.0,0.0,0.0}-\mathrm{div}\varepsilon\mathrm{grad}&0&\color[rgb]{0.0,0.0,0.0}-\mathrm{div}\varepsilon\mathrm{grad}\\ 0&0&\mu&0\\ \color[rgb]{0.0,0.0,0.0}-\mathrm{div}\varepsilon&\color[rgb]{0.0,0.0,0.0}-\mathrm{div}\varepsilon\mathrm{grad}&0&\color[rgb]{0.0,0.0,0.0}-\mathrm{div}\varepsilon\mathrm{grad}\end{array}\right]}_{{\mathbf{E}}_{\mathrm{EMQS4}}}\frac{\partial}{\partial t}\left[\begin{array}[]{c}\!\frac{\partial}{\partial t}\bm{A}\!\\ \!\varphi\!\\ \!\bm{H}\!\\ \!\lambda\!\end{array}\right]\!\!=\!\!\ \left[\left[\begin{array}[]{cccc}0&0&-\mathrm{curl}&0\\ 0&0&0&0\\ \mathrm{curl}&0&0&0\\ 0&0&0&0\end{array}\right]\right.\\ -\left.\left[\begin{array}[]{cccc}\kappa&\kappa\mathrm{grad}&0&0\\ \color[rgb]{0.0,0.0,0.0}-\mathrm{div}\kappa&\color[rgb]{0.0,0.0,0.0}-\mathrm{div}\kappa\mathrm{grad}&0&0\\ 0&0&0&0\\ 0&0&0&0\end{array}\right]\right]\left[\begin{array}[]{c}\!\frac{\partial}{\partial t}\bm{A}\!\\ \!\varphi\!\\ \!\bm{H}\!\\ \!\lambda\!\end{array}\right]\!\!+\!\!\left[\begin{array}[]{c}\!1\!\\ \!\color[rgb]{0.0,0.0,0.0}-\mathrm{div}\!\\ \!0\!\\ \!\color[rgb]{0.0,0.0,0.0}-\mathrm{div}\!\end{array}\right]\bm{j}_{\mathrm{s}},\end{split}$ (16)
@@
$H_{\mathrm{EMQS}}=\int_{\Omega}\left\{\color[rgb]{0,0,0}\frac{1}{2}\left(\varepsilon\|\mathrm{grad}\varphi\|^{2}+\varepsilon\|\mathrm{grad}\lambda\|^{2}+\mu\|\bm{H}\|^{2}\right)\right.\\ \left.+\varepsilon\frac{\partial}{\partial t}\bm{A}\cdot\mathrm{grad}(\varphi+\lambda)+\varepsilon\mathrm{grad}\varphi\cdot\mathrm{grad}\lambda\right\}d\Omega$ (17)
corresponding the EMQS energy, (4) holds for the formulation (16) with
$\mathrm{grad}\,H_{\mathrm{EMQS}}=\mathbf{E}_{\mathrm{EMQS4}}^{\star}{\mathbf{z}}=\left[\begin{array}[]{c}\varepsilon\mathrm{grad}(\varphi+\lambda),\\ \color[rgb]{0.0,0.0,0.0}-\mathrm{div}\left(\varepsilon(\frac{\partial}{\partial t}\bm{A}+\mathrm{grad}(\varphi+\lambda))\right)\\ \mu\bm{H}\\ \color[rgb]{0.0,0.0,0.0}-\mathrm{div}\left(\varepsilon(\frac{\partial}{\partial t}\bm{A}+\mathrm{grad}(\varphi+\lambda))\right)\end{array}\right].$
@@
Symmetrized EMQS field formulations can also be set up to explicitly enforce the Coulomb-type gauge condition (12) via $\mathrm{div}\left(\hat{\kappa}\frac{\partial}{\partial t}\bm{A}\right)=\frac{1}{\tau}\mathrm{div}\left(\varepsilon\frac{\partial}{\partial t}\bm{A}\right)=0,$ with an artificial $\hat{\kappa}:=\frac{1}{\tau}\varepsilon$ (for a time constant $\tau$ to get the correct units) as
$\begin{split}\underbrace{\left[\begin{array}[]{cccc}0&\varepsilon\mathrm{grad}&0&0\\ \color[rgb]{0.0,0.0,0.0}-\mathrm{div}\varepsilon&\color[rgb]{0.0,0.0,0.0}-\mathrm{div}\varepsilon\mathrm{grad}&0&0\\ 0&0&\mu&0\\ 0&0&0&0\end{array}\right]}_{{\mathbf{E}}_{\mathrm{EMQS5}}}\frac{\partial}{\partial t}\left[\begin{array}[]{c}\!\frac{\partial}{\partial t}\bm{A}\!\\ \!\varphi\!\\ \!\bm{H}\!\\ \!\lambda\!\end{array}\right]\!\!=\!\!\ \left[\left[\begin{array}[]{cccc}0&0&-\mathrm{curl}&{-\hat{\kappa}\mathrm{grad}}\\ 0&0&0&0\\ \mathrm{curl}&0&0&0\\ {-\mathrm{div}\hat{\kappa}}&0&0&0\end{array}\right]\right.\\ -\left.\left[\begin{array}[]{cccc}\kappa&\kappa\mathrm{grad}&0&0\\ \color[rgb]{0.0,0.0,0.0}-\mathrm{div}\kappa&\color[rgb]{0.0,0.0,0.0}-\mathrm{div}\kappa\mathrm{grad}&0&0\\ 0&0&0&0\\ 0&0&0&0\end{array}\right]\right]\left[\begin{array}[]{c}\!\frac{\partial}{\partial t}\bm{A}\!\\ \!\varphi\!\\ \!\bm{H}\!\\ \!\lambda\!\end{array}\right]\!\!+\!\!\left[\begin{array}[]{c}\!1\!\\ \!\color[rgb]{0.0,0.0,0.0}-\mathrm{div}\!\\ \!0\!\\ \!0\!\end{array}\right]\bm{j}_{\mathrm{s}},\end{split}$ (19)
with the Lagrange multiplier $\lambda$, input $\mathrm{u}(t)=\bm{j}_{\mathrm{s}}$ and $\mathrm{y}(t)=\left(\frac{\partial}{\partial t}\bm{A}+\mathrm{grad}\varphi\right)=-\bm{E}.$ Note that $\hat{\kappa}$ can also be defined domain dependent, i.e., the gauge (12) can be defined in the whole computational domain or just in the non-conducting domains with the original electric conductivity $\kappa=0.$ The system (19) is a PHS, where the compatibility relation (4) holds for the Hamiltonian (14). It maintains stability and energy conservation also for the case $\kappa=0$, i.e., in the absence of ohmic losses. With the explicitly enforced Coulomb-type gauge (12), the Hamiltonian (14) of the symmetric formulation (15) reduces to the expression $H_{\mathrm{EMQSR}}=\frac{1}{2}\left(\varepsilon\|\mathrm{grad}\varphi\|^{2}+\mu\|\bm{H}\|^{2}\right).$ This expression corresponds the energy of the EMQS field approximation, i.e., a formulation with the combined magnetic field and electro-quasistatic field energy.
@@ 4 Conclusions
Various infinite dimensional EMQS formulations have been classified within the port-Hamiltonian system framework, thus implying their numerical stability after a proper mimetic discretization where the PHS structure is found. The validity of the compatibility condition could be shown for symmetrized formulations, i.e., those EMQS formulations based on the Darwin-Ampere and the full Maxwell continuity equation. However, this re-introduced the second order time derivative of the magnetic vector potential that was discarded in the Darwin-Ampere equation under the no radiation EMQS assumption. For explicitely gauged EMQS formulations the relation between Lagrange multiplier and split scalar potential formulations is shown.